OpenStudy (anonymous):

The perimeter of a rectangle is the sum of all of its sides. Write a mathematical expression for the perimeter of a rectangle with a length of 4 and a width of w. Simplify the expression if possible.

OpenStudy (anonymous):

perimeter of rectangle: 2L + 2W given L=4, W=w we have perimeter: 2(4) + 2w = 8 + 2w or 2(4+w)